OpenStudy (nathaliama13):

pls help !!! Look at the graph below: First line joins ordered pairs negative 4, 3 and 2, negative 3. Second line joins negative 4, negative 3 and 2, 3. Part A shaded above first and second line. Part B shaded below first line and above second line. Part C shaded below first and second lines. Part D shaded above first line and below second line. Which part of the graph best represents the solution set to the system of inequalities y ≥ x + 1 and y + x ≥ −1? Part A Part B Part C Part D

OpenStudy (anonymous):

You're better off drawing this for us. It's pretty tough to translate that text into a graph.

OpenStudy (anonymous):

Well, rewrite y + x >= -1 similar to how y>= x + 1 is written. Then y >= -x - 1 is the other boundary. HINT: The black lines separating the different regions are the equations: y = x + 1 and y = -x - 1. So tell me, If I have to be greater than y = x + 1, would I be above or below the line going from the bottom left to the upper right?

OpenStudy (nathaliama13):

you would be above, i think? i'm not so good at inequalities :(

OpenStudy (nathaliama13):

@adam.bloedel hello ?

OpenStudy (anonymous):

Sorry. I had to go to work yesterday. Indeed you would be. So you would have to be in either the yellow or red region. But you also know that you are above y = -x - 1 which is the line running from top left to the bottom right. Knowing this... which region would be above both lines?